Posted: 6/18/2013 9:24:19 AM
Louisiana Business Emergency Operations Center


022-13 Pipeline Leak Franklin Parish (Delhi) 13 June 2013





At 1402 hours 13 June 2013, a notification was received of a
pipeline leak in Franklin Parish near the community of Delhi.  Denbury
Resources and Trans Canada are the two companies that have pipelines in the
area.  Denbury shut in all of their lines and confirmed the leak was not
theirs.  Trans Canada confirmed the leak was in their lines and that the
product leaking is methane gas. 


17 June 2013

At approximately 1400 hours,
personnel discovered an area approximately one half mile from the original
release location where another release had begun.  This second release was
not in existence this morning.  This second release is gurgling up in the
swamp in several places and elevating the CO2 and LEL levels in this
area.  At 1430 hours another meeting was held with all involved
parties.  The original release has almost completely stopped and it is
believed that the trapped gases have found another path to the surface
resulting in the second release.  Crews have been reassigned to contain
the by-products of the release and the plan for containment of the area seems
adequate.  It is estimated the second release to be between 2,000 and
2,500 barrels a day.  There is no additional off site impact at this
time.  Air monitoring continues to be in place


As of 0943 hours, the incident was
ongoing. Boots and Coots had made it to within 60 yards of the source. Air
monitoring is ongoing with no increase in readings. Pumping operations to
bridge the source is ongoing with progress being made.


16 June 2013:

At 0507 hours it was reported that workers were backed up
from the collection site due to a raised VOC and CO2 reading from air monitors.
Collection will continue once reading drop again. Boots and Coots continue to
pump heavy water and mud into the system to try to stop the leak. As of 1046
hours, the readings were still too high to approach the well site.


15 June 2013

As of 0822 hours, it was reported that the incident action
plan for 15 June was finished. Continuous air monitoring was ongoing. Readings
taken at all evacuated areas were good. The evacuation is still in effect but
was being evaluated. The road closure for Antely Road was still in place. The
evacuation order was lifted at 1400 hours with each evacuee to be escorted to
their homes by a Sheriff’s Deputy and a representative from CTEH to conduct air
quality testing at each residence to ensure the residence was safe. Boots and
Coots began dumping heavy water into the system to try and shut the well in.
Aerial surveillance showed the source was still leaking what appeared to be
drilling mud which was now flowing and not cascading as it was on 14 June.


As of 2000 hours, a containment damn was installed.


14 June 2013

As of 1634 hours 14 June, it was reported that Denbury is
actually the responsible party. A well was blown out and Boots and Coots were
engaged to plug the well with the work planned for 15 June. Evacuations were
being reevaluated.  As of 1857 hours, the incident was ongoing with
evacuations still in place. Air monitoring was ongoing by USES and CTEH.
Indications were that this incident would continue for several days. LSP, DEQ,
Richland Parish Sheriff’s Office and both Richland and Franklin OEP directors
were on site.


As of 0800 hours 14 June, the wrong line was blown the night
of 13 June and the leak is ongoing and pressure readings on the line are not
dropping. Trans Canada has a contractor en route. Antely Road remains closed
and the evacuation is still in effect.


13 June 2013

As of 1552 hours 13 June, LSP Hazmat reported that
monitoring and identification of the leaking line was ongoing and La Hwy 132
remained closed as a precautionary measure.


There have been no reports of any explosions or fires
associated with the leak.


There have been no injuries or fatalities reported for this


Initially, 7 homes were evacuated as a precautionary measure
and one road (Antley Road in Franklin Parish) was closed. As of 1630 hours, 5
additional homes were evacuated. Residents are being set up in a local hotel by
Trans Canada. The evacuation order was lifted  at 1400 hours 15 June with
residents escorted back home and their residences tested for air quality by


No requests for state assistance have been submitted.



OEP Directors for both Franklin and Richland Parish are




GOHSEP Region 8 Coordinator is assisting the parish affected
and providing reports as needed.


WebEOC incident 022-13 Pipeline Leak Franklin Parish (Delhi)
13 June 2013 has been created to support this event.


GOHSEP Crisis Action Team was activated at 1500 hours 14
June 2013 and will remain activated until further notice.



Personnel responded to the incident site to monitor and
report as needed and will maintain personnel on site until secured.


Conducting aerial reconnaissance to assist LSP Emergency
Services personnel on with the assessment of the leak.




Personnel are performing air monitoring and will continue to
do so until the site is secure


GOHSEP Crisis Action Team will continue to monitor, report
and assist with this incident as needed.



The LA BEOC Portal is designed and developed by the NIMSAT Institute under the Cooperative Endeavor Agreement with the Louisiana Governor's Office of Homeland Security and Emergency Preparedness (GOHSEP)


© Copyright 2010-2019, NIMSAT. All rights reserved. Terms of Service. Portal Guidelines.